Nippons Steels expenses for heating and cooling a large manufacturing facility are expected to increase according to an arithmetic gradient beginning in year 2. If the cost is $550,000 this year (year 0) and will be $550,000 again in year 1, but then are estimated to increase by $46,000 each year through year 12, what is the equivalent annual worth in years 1 to 12 of these energy costs at an interest rate of 14% per year?
Answer #1

First we estimate the PW of future expenses

PW=550000+550000*(P/A,0.14,12)+46000*(P/G,0.14,12)

PW=550000+550000*5.660292+46000*22.639878=$4704594.99

EUAW=PW/(P/A,0.14,12)=4704594.99/5.660292=$831,157.65